# Agent Journey Beyond RGB: Unveiling Hybrid Semantic-Spatial Environmental Representations for Vision-and-Language Navigation

### v4

@@ Abstract
Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN). Existing approaches primarily rely on RGB images for environmental representation, underutilizing latent textual semantic and spatial cues and leaving the modality gap between instructions and scarce environmental representations unresolved. Intuitively, humans inherently ground semantic knowledge within spatial layouts during indoor navigation. Inspired by this, we propose a versatile Semantic Understanding and Spatial Awareness (SUSA) architecture to encourage agents to ground environment from diverse perspectives. SUSA includes a Textual Semantic Understanding (TSU) module, which narrows the modality gap between instructions and environments by generating and associating the descriptions of environmental landmarks in agent’s immediate surroundings. Additionally, a Depth-enhanced Spatial Perception (DSP) module incrementally constructs a depth exploration map, enabling a more nuanced comprehension of environmental layouts. Experiments demonstrate that SUSA’s hybrid semantic-spatial representations effectively enhance navigation performance, setting new state-of-the-art performance across three VLN benchmarks (REVERIE, R2R, and SOON). The source code will be publicly available.
@@ 1 Introduction
With the advancement of multimodal technology [44, 47, 29, 59], several embodied VLN tasks [2, 41, 37, 8] have garnered significant attention due to their promising applications in robotics and intelligent assistance. Among them, conventional VLN tasks [2, 27] focused on step-by-step navigation within unseen environments based on detailed instructions, while goal-oriented VLN tasks [64, 41, 45] demand agents to identify predefined objects based on high-level instructions. Consequently, VLN agents are required to profoundly cognize the spatial information of environment and ground them with textual instructions to brilliantly accomplish arduous navigation tasks.
@@
Figure 1: An overview of the proposed SUSA. Unlike previous approaches that solely rely on RGB input, we introduce view-level textual panoramas and trajectory-level depth exploration maps, supporting the agent’s explicit understanding of the environment.
Considerable literature has exhibited the efficacy of Transformer-based models [5, 20, 38, 52, 40] for interacting with instructions, RGB-based environmental observations, and historical trajectories. However, we suggest that RGB-only visual environment representations leave VLN fraught with two challenges: the modality gap between instructions and the environment, and the scarcity of training environmental representations. Firstly, VLN agents struggle to precisely match the landmarks in language instructions with the entities in visual environments, due to the intrinsic modality heterogeneity. To address this, recent works [12, 32] utilized contrastive learning to promote alignment, while others leveraged candidate object concepts [50, 33], large language models (LLMs) [34] or retrieved knowledge [31] to explicitly reason the textual semantics (i.e., linguistic descriptions about landmarks) to enhance environmental representations. However, these methods do not directly align textual semantics with landmarks (e.g., “fireplace” in Fig. 1) in instruction, yielding trivial gain.
The second challenge pertains to the monotonous and limited environmental representations, which often leads agents to overfit to familiar spatial layouts [51]. To alleviate this limitation, recent endeavors [55, 7] augment the quantity of environmental scenes, while others diversify environmental representations by constructing graphs [16, 23], grids [56, 1], or volumetric environments [36]. However, these RGB-based methods easily suffer from attention imbalance or environmental bias [65, 61]. FDA [18] utilizes frequency information to enrich environmental representations. However, depth images vividly convey spatial information (e.g., doors and corridors), which is often overlooked in discrete navigation tasks [49]. Few works [1, 56] incorporate depth, they merely project depth values onto the corresponding spatial locations of RGB images, rather than independently grounding depth[53]. Overall, the underlying textual semantic and spatial representations remains underutilized, though studies [22] suggest that injecting inductive biases where different modules access distinct modalities, offers a promising solution.
In this work, we propose the Semantic Understanding and Spatial Awareness (SUSA) architecture, which individually grounds semantic and spatial representations of the environment with instructions and synergistically predicts navigation actions from diverse perspectives. As illustrated in Fig. 1, we first construct a view-level textual panorama by extracting salient landmarks as the textual semantics to explicitly match them with language instructions, and select the most instruction-relevant view, thereby bridging the semantic gap between the visual and linguistic modalities. Further, beyond RGB images, we progressively construct an independent trajectory-level depth exploration map to enhance the agent’s spatial perception of the environment, also directly aligning it with instructions. To enable the agent to absorb complementary information from different environmental perspectives, we aggregate these hybrid semantic-spatial environmental representations to synergistically facilitate hierarchical navigation predictions, and employ contrastive learning to align these hybrid representations with language instructions. We empirically substantiate that exclude semantic panoramas and depth exploration maps during pretraining manifest markedly enhanced generalization capabilities. Experimental results demonstrate that our agent achieves significantly improved navigation performance, outperforming state-of-the-art methods across three vision-and-language navigation benchmarks: R2R [2], REVERIE [41], and SOON [64].
The main contributions of our SUSA are as follows:
@@
  • SUSA devises a textual-aware semantic understanding (TSU) module to match the most related landmarks between environments and instructions, explicitly addressing the modality heterogeneity.
@@
  • SUSA further presents a depth-enhanced spatial perception (DSP) module to explore and represent environmental layouts beyond surface RGB content.
  • Extensive experiments on three VLN benchmarks have demonstrated the superiority of hybrid semantic-spatial environmental representations introduced by SUSA.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). Vision-and-Language Navigation (VLN) tasks [2, 41, 64] require an agent to navigate in embodied environments sequentially towards a designated location or object based on natural language instructions. Given that the VLN task can be formulated as a partially observable Markov decision process, early methods [54, 66, 14, 26] primarily relied on reinforcement learning and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss [63, 54] designs. More recently, to better generalize to unknown environments and ground instructions, many researchers [5, 20, 38, 52, 40] have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories. While these approaches have propelled VLN forward, they mainly rely on fragile RGB-based environment representations. This paper encourages the agent to perceive and ground the environment from text semantics and depth perspectives.
Semantic Alignment in VLN. To align the semantic relationship between the environment and instructions, current practitioners in the VLN community can be divided into two categories: implicit and explicit approaches. Implicit methods [12, 51, 32] commonly employ contrastive learning or attention mechanisms to associate visual contexts with textual instructions. However, given the inherent semantic gap between these modalities, implicit approaches have limited alignment capabilities. Explicit alignment techniques aim to project both semantics into a same modality space. For example, some methods transform the environment into a text-based modality, Console [34] scores landmark co-occurrences to identify key objects, and AACL [33] leverages CLIP [44] to refine textual entities. Similar to our approach, KERM [31] utilizes retrieved knowledge as textual semantics for each view. However, these explicit methods directly “stitch” textual and visual environment representations, which may lead to an overreliance on text modality yet underutilizing visual data [67]. We thus take environmental textual semantic as an independent modality to directly match the most instruction-relevant landmarks.
@@
Environment Representations in VLN. Limited training sources and environmental representations in VLN tasks hinder the agent to generalize the unknown layouts. Prior endeavors mitigate this limitation through data augmentation [35, 55] or by diversifying visual representations [60, 1, 56, 36]. However, these approaches primarily rely on RGB images, while alternative modalities such as frequency [18], spectral[24], depth [49] and semantic information [53, 21, 4] remain relatively underexplored. Among them, SEAT [53] enhances cross-domain environmental perception by querying the Matterport3D simulator[3] for depth and semantics. However, SEAT fuses these multi-type representations into a unified space without explicitly grounding each modality. Moreover, these semantic representations [53, 21] are not inherently textual, limiting their ability to bridge the modality gap. In contrast, our method constructs the view-level textual panoramas and trajectory-level depth exploration maps, directly grounding them with instructions to perceive rich environmental representations.
@@
Figure 2: The detailed architecture of our proposed SUSA model. The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively. The Hybrid Representation Aggregation (HRA) module is designed for aligning environmental representations with instructions.
@@ 3.1 Preliminary
Task Formulation. We focus on discrete VLN tasks [2, 41, 64] within a Matterport3D simulator [3], modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths. At the outset of each episode, the agent equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words. We use BERT [10] to extract instruction features and then embed them as =$\mathcal{I}=\{w_{1},\dotso,w_{l}\}$. At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views. During navigation, the agent make the next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$. The sequential decision-making process is iterated until the agent decides to stop at the current location or exceeds the maximum navigation steps. For goal-oriented tasks [41, 64], the agent is also tasked with locating a target object.
Overview of the Proposed Method. Our objective is to accord complementary environmental content by equipping the agent with a textual semantic panorama and a depth exploration map. Following prevailing approaches [18, 30, 19], we adopt DUET [6] as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions. The SUSA framework consists of three main modules: (a) textual-aware semantic selection is detailed in Sec. 3.2, (b) depth-aware visual environment exploration is presented in Sec. 3.3, and (c) hybrid representation aggregation and hierarchical action prediction is subsequently introduced in Sec. 3.4, as illustrated in Fig. 2.
@@ 3.2 Textual-Aware Semantic Understanding
Figure 3: Illustration of TSU module in  Fig. 2, which matches instructions and semantic features by static and dynamic strategies.
Salient Landmarks Extractor. The pretrained visual-language model readily provides the capability for agents to perceive textual semantics by generating detailed captions from RGB images. As illustrated in Fig. 2, we utilize off-the-shelf BLIP-2-FlanT5-xxL [29], a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”. Following [15, 28], this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks. These descriptions typically involve room types and objects (e.g.,“a living room with two chairs”), which assist in the object recognition. We further discuss the quality of the generated descriptions in supplementary materials. Additionally, we employ the CLIP (ViT-L-14-336px) text encoder [44] to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$. Thanks to the powerful yet frozen BLIP-2 and CLIP models, we empower the agent to obtain textual semantic features from the navigable nodes in its vicinity during navigation.
Textual Semantic Selector. Previous methods [50, 33, 31] aligned environmental RGB images with instructions by exploiting textual semantics yet the inherent modality discrepancy between visual and language modalities remained unaddressed. Similar to [39], we treat the textual semantic as a perceptual representation for navigation in this module, directly matching salient landmarks with the instructions. As shown in Fig. 2 and Fig. 3, our textual semantic selector combines both $static$ and $dynamic$ matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
@@
$\small\begin{split}\mathbf{M}_{i,j}=\frac{\mathcal{X}^{\mathcal{S}}_{t,i}\cdot\mathcal{I}_{j}}{\|\mathcal{X}^{\mathcal{S}}_{t,i}\|\|\mathcal{I}_{j}\|},\ \ \ \ \Gamma_{t}^{stat}=\max_{j}\mathbf{M}_{i,j}\end{split}$ (1)
@@
$\small\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{X}^{\mathcal{S}}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
@@ 3.3 Depth-Enhanced Spatial Perception
Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited in earlier works. To enhance the spatial perception, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one [6].
Depth-Enhanced Exploration Mapping. Concretely, we obtain ground truth depth images for each view from the Matterport3D simulator. As illustrated in Fig. 2, we then employ ResNet-50 [17] (pretrained on the Gibson [57]) and CLIP (ViT-L/14@336px) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view. Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images. The panorama encoding modules for both modalities share weights, since each view in the depth and RGB panoramas is accordant. Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel. The explored depth and RGB trajectory maps are formalized as $\mathcal{T}_{t}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node. These exploration maps memorize the agent navigation trajectory information for global action predictions, which serves to guide the agent traversal across various navigable nodes when considering backtracking.
Cross-Modal Interaction and Reasoning. To defer to instructions, we utilize a multi-layer cross-encoder that interacts with both the language instructions and the depth exploration map features to generate the refined depth map $\tilde{\mathcal{T}^{\mathcal{D}}}$. This encoder comprises four residual connection layers of a cross-modal transformer, each incorporating graph-aware self-attention (GASA) [6], cross-modal attention, and a feed-forward network. The GASA operation is formally expressed as $\text{GASA}(X)=\eth(X\Theta_{q}(X\Theta_{k})^{T}+\mathcal{M})X\Theta_{v}$ where $\mathcal{M}$ represents a learnable distance matrix encoding the relative distances between navigation nodes, $\eth$ denotes the Softmax activation function, and $\Theta_{q},\Theta_{k},\Theta_{v}$ are learnable weight matrices. In parallel, we facilitate similar cross-modal interactions between the language instructions and the RGB exploration map $\mathcal{T}^{\mathcal{R}}$, as well as the RGB panoramas $\mathcal{X}^{\mathcal{R}}$. These are encoded as $\tilde{\mathcal{T}^{\mathcal{R}}}$ and $\tilde{\mathcal{X}^{\mathcal{R}}}$, thereby enabling the model to perform both global and local predictions based on the given instructions $\mathcal{I}$. Compared to the standalone $\mathcal{T}^{\mathcal{R}}$, the incorporation of a depth exploration map $\mathcal{T}^{\mathcal{D}}$ enhances the perceptual clarity of environmental structures, mitigating the risk of overfitting to RGB-based visual noises [25] or biases [22, 51].
Figure 4: Illustration of the pipeline for fusing hybrid environmental representations in the proposed HRA module in Fig. 2.
3.4 Hybrid Representation Aggregation

### v5

# Agent Journey Beyond RGB: Hierarchical Semantic-Spatial Representation Enrichment for Vision-and-Language Navigation

@@ Abstract
Navigating unseen environments based on natural language instructions remains difficult for egocentric agents in Vision-and-Language Navigation (VLN). Intuitively, humans inherently ground concrete semantic knowledge within spatial layouts during indoor navigation. Although previous studies have introduced diverse environmental representations to enhance reasoning, other co-occurrence modalities are often naively concatenated with RGB features, resulting in suboptimal utilization of each modality’s distinct contribution. Inspired by this, we propose a hierarchical Semantic Understanding and Spatial Awareness (SUSA) architecture to enable agents to perceive and ground environments at diverse scales. Specifically, the Textual Semantic Understanding (TSU) module supports local action prediction by generating view-level descriptions, thereby capturing fine-grained environmental semantics and narrowing the modality gap between instructions and environments. Complementarily, the Depth-enhanced Spatial Perception (DSP) module incrementally constructs a trajectory-level depth exploration map, providing the agent with a coarse-grained comprehension of the global spatial layout. Extensive experiments demonstrate that SUSA’s hierarchical representation enrichment not only boosts the navigation performance of the baseline on discrete VLN benchmarks (REVERIE, R2R, and SOON), but also exhibits superior generalization to the continuous R2R-CE.
Codehttps://github.com/HCI-LMC/VLN-SUSA
@@ 1 Introduction
With the advancement of multimodal technology (Radford et al. 2021; Li et al. 2023a; Zhang et al. 2024) in the artificial intelligence community, embodied VLN tasks (Krantz et al. 2020; Qi et al. 2020; Das et al. 2018) have garnered significant attention due to their promising applications in robotics and intelligent assistance. Conventional VLN tasks (Anderson et al. 2018; Krantz et al. 2020) focused on step-by-step navigation in unseen environments based on detailed instructions, while goal-oriented VLN tasks (Zhu et al. 2021; Qi et al. 2020) demand agents to identify predefined objects based on high-level instructions. Consequently, VLN agents are required to profoundly cognize the environmental spatial information and ground them with textual instructions to brilliantly accomplish navigation tasks.
@@
Figure 1: An overview of the proposed SUSA. Beyond RGB inputs, we introduce view-level textual panoramas and trajectory-level depth exploration maps, supporting the agent’s explicit understanding of the environment.
Recently, increasing the diversity and quantity of augmented environments (Wang et al. 2023b, c; Liu, Wang, and Yang 2024) has emerged as a popular strategy to facilitate the generalization ability of VLN agents in unseen scenarios. This trend is largely motivated by the limited navigational environments in simulators, as well as the powerful fitting capacity of Transformer-based VLN models (Chen et al. 2021; Pashevich, Schmid, and Sun 2021). Nonetheless, VLN agents still struggle to accurately ground the landmarks mentioned in language instructions with the RGB-based visual scene, due to the intrinsic modality heterogeneity. To address this, recent studies introduce textual semantics as perceptual representations to guide navigation (Pan et al. 2024). These approaches leverage candidate object concepts (Wang et al. 2023a; Lin et al. 2023), large language models (LLMs) (Lin et al. 2024), or external knowledge (Li et al. 2023b) to explicitly reason about textual descriptions of landmarks and enhance environmental grounding. However, these methods tend to combine textual and visual representations in a straightforward manner, which limits the effective utilization of each modality and constrains the interpretability of their individual contributions. Similarly, co-occurring environmental cues from other modalities—such as frequency (He et al. 2024), spectrum (Hwang et al. 2023), and depth (Tan et al. 2022)—have also been preliminarily explored, though remain underutilized.
In this work, we propose the hierarchical Semantic Understanding and Spatial Awareness (SUSA) architecture, which augments the agent’s environmental comprehension by effectively perceiving and exploiting fine-grained textual semantics and coarse-grained depth-aware spatial information. To enable fine-grained semantic grounding, as shown in Fig. 1, we construct a view-level textual panorama by extracting salient landmarks as textual semantics, thereby bridging the gap between visual and linguistic modalities. The most instruction-relevant view is then selected for local navigation prediction. For global action prediction, we progressively construct a trajectory-level depth exploration map to strengthen the agent’s coarse-grained spatial perception. Finally, SUSA hierarchically aggregates these hybrid semantic-spatial representations to perceive complementary environmental information and employs contrastive learning to align them with natural language instructions. Extensive experiments on four VLN benchmarks (R2R (Anderson et al. 2018), REVERIE (Qi et al. 2020), SOON (Zhu et al. 2021) and R2R-CE (Krantz et al. 2020)) demonstrated the superiority of hybrid semantic-spatial representations enrichment. The contributions of our SUSA are as follows:
@@
  • SUSA devises a textual-aware semantic understanding (TSU) module to perceive fine-grained textual semantics of environments, and explicitly address modality heterogeneity between instructions and environments.
@@
  • SUSA further constructs a depth exploration map within the depth-enhanced spatial perception (DSP) module to complementarily perceive global environmental layouts and reinforce holistic spatial awareness beyond RGB.
  • Experiments demonstrate that SUSA substantially improves upon the baseline, achieving state-of-the-art performance on three discrete VLN tasks, and exhibits promising generalization in continuous environments.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). Vision-and-Language Navigation (VLN) tasks (Anderson et al. 2018; Qi et al. 2020; Zhu et al. 2021) require an agent to navigate in embodied environments sequentially towards a designated location or object based on natural language instructions. Given that the VLN task can be formulated as a partially observable Markov decision process, early methods (Wang et al. 2019; Fried et al. 2018) primarily relied on reinforcement learning and imitation learning paradigms to improve navigation performance, incorporating various strategies and auxiliary loss (Wang et al. 2019) designs. More recently, to better model the relationship between the environment and instructions, many researchers (Chen et al. 2021; Pashevich, Schmid, and Sun 2021) have focused on transformer-based architectures, which effectively model interactions among instructions, environmental features, and historical navigation trajectories.
@@
Environment Representations in VLN. Limited training sources and environmental representations in VLN tasks hinder the agent to generalize the unknown layouts. To alleviate this limitation, recent endeavors (Wang et al. 2023b; Chen et al. 2022a) augment the quantity of environmental scenes, while others diversify environmental representations by constructing graphs (Georgakis et al. 2022; Huang et al. 2023), grids (Wang et al. 2023c; An et al. 2023), or volumetric environments (Liu, Wang, and Yang 2024). Beyond RGB inputs, recent efforts have explored other co-occurrence multimodal(Li et al. 2023b; He et al. 2024) cues to enrich environmental representations. For instance, SEAT (Wang et al. 2024b) enhances cross-domain environmental perception by querying the Matterport3D simulator (Chang et al. 2017) for depth and semantics. However, SEAT fuses these multi-type representations into a unified space without explicitly grounding each modality. Moreover, these semantic representations (Wang et al. 2024b; Hong et al. 2023) are not inherently textual, limiting their ability to bridge the modality gap. Other approaches incorporate depth information (An et al. 2023; Wang et al. 2023c); however, they typically project depth values onto the spatial coordinates of RGB images without independently grounding the depth modality. In contrast, our method constructs the view-level textual panoramas and trajectory-level depth exploration maps, individually grounding them with instructions to perceive rich environmental representations.
@@
Figure 2: The detailed architecture of our proposed SUSA model. The orange and green arrows highlight our proposed Textual-Aware Semantic Understanding (TSU) and Depth-Aware Spatial Perception (DSP) modules, respectively. The Hierarchical Aggregation and Prediction (HAP) module is designed for hierarchical aligning environmental representations with instructions.
@@ 3.1 Preliminary
Task Formulation. We focus on discrete VLN tasks (Anderson et al. 2018; Qi et al. 2020; Zhu et al. 2021) within a Matterport3D simulator (Chang et al. 2017), modeled as an undirected graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, where $\mathcal{N}$ denotes nodes and $\mathcal{E}$ signifies the connectivity paths. At the outset of each episode, the agent equipped with a GPS sensor, RGB and depth cameras, is situated at a randomly navigable node and receives a language instruction $I=\{\text{w}_{i}\}_{i=1}^{l}$ including $l$ words. We use BERT (Devlin et al. 2019) to extract instruction features and then embed them as $\mathcal{I}=\{w_{1},\dotso,w_{l}\}$. At each time step $t$, the agent can perceive the surrounding RGB panorama $R_{t}=\{\text{r}_{t,i}\}_{i=1}^{n=36}$ and depth panorama $D_{t}=\{\text{d}_{t,i}\}_{i=1}^{n=36}$ along with $n$ corresponding views. During navigation, the agent makes the next action $a_{t}$ by selecting one navigable node from the candidate actions $\mathcal{A}_{t}=\{a_{t,i}\}_{i=0}^{k}$.
Overview of the Proposed Method. Our objective is to enrich complementary environmental content by equipping the agent with a textual semantic panorama and a depth exploration map. Following prevailing approaches (He et al. 2024), we adopt DUET (Chen et al. 2022b) as our baseline, which merely utilizes RGB panoramas and topological maps as environmental representations for both local and global action predictions. The SUSA framework consists of three main modules: (a) TSU: textual-aware semantic selection is detailed in Section 3.2, (b) DSP: depth-enhanced visual environment exploration is presented in Section 3.3, and (c) HAP: hybrid representation aggregation and hierarchical action prediction is subsequently introduced in Section 3.4, as illustrated in Fig. 2.
@@ 3.2 Textual-Aware Semantic Understanding
Figure 3: Illustration of the TSU module in Fig.2, which matches instructions and semantic features by static and dynamic matching strategies.
Salient Landmarks Extractor. The pretrained visual-language model readily provides the capability for agents to perceive textual semantics by generating detailed captions from RGB images. As illustrated in Fig. 2, we utilize off-the-shelf BLIP-2-FlanT5-xxL (Li et al. 2023a), a powerful visual-language model, to discretize $R_{t}$ and generate corresponding textual descriptions for each view using the simple prompt “a photo of”. Following (Gao et al. 2024; Li and Bansal 2023), this process yields a textual panorama $\mathcal{S}_{t}=\{s_{t,i}\}_{i=1}^{n=36}$, where each view $s_{t,i}$ contains informed descriptions of salient landmarks. These descriptions typically involve room types and objects (e.g.,“a living room with two chairs”), which assist in the object recognition. We further discuss the quality of the generated descriptions in supplementary materials. Additionally, we employ the CLIP (ViT-L-14-336px) text encoder (Radford et al. 2021) to extract textual semantic panorama features, embedding them as $\mathcal{X}^{\mathcal{S}}_{t}=\{x^{s}_{t,1},\dots,x^{s}_{t,n}\}$. Thanks to the powerful yet frozen BLIP-2 and CLIP models, we empower the agent to obtain textual semantic features from the navigable nodes in its vicinity during navigation.
Textual Semantic Selector. Similar to (Pan et al. 2024), we treat the textual semantic as a perceptual representation for navigation in this module, directly matching salient landmarks with the instructions. As shown in Fig. 2 and Fig. 3, our textual semantic selector combines both staticstatic and dynamicdynamic matching strategies to select the most instruction-relevant view in the textual panorama $\mathcal{X}^{\mathcal{S}}_{t}$.
@@
$\begin{split}\mathbf{M}_{i,j}=\frac{\mathcal{X}^{\mathcal{S}}_{t,i}\cdot\mathcal{I}_{j}}{\|\mathcal{X}^{\mathcal{S}}_{t,i}\|\|\mathcal{I}_{j}\|},\ \ \ \ \Gamma_{t}^{stat}=\max_{j}\mathbf{M}_{i,j}\end{split}$ (1)
@@
$\begin{split}\Gamma_{t}^{dyn}=\text{TCA}(\mathcal{X}^{\mathcal{S}}_{t},\mathcal{I},\mathcal{I}),\end{split}$ (2)
@@ 3.3 Depth-Enhanced Spatial Perception
Depth images, which also contain more intuitive and easily distinguishable structural information, were not adequately exploited in earlier works (Tan et al. 2022). To enhance the spatial perception, we concurrently construct a depth-based exploration map to memorize depth-based navigation trajectory alongside the RGB-based one (Chen et al. 2022b).
Depth-Enhanced Exploration Mapping. Concretely, we obtain ground truth depth images for each view from the Matterport3D simulator. As illustrated in Fig. 2, we then employ ResNet-50 (He et al. 2016) (pretrained on the Gibson (Xia et al. 2018)) and CLIP (ViT-L/14@336px) as the depth and RGB encoders, respectively, to extract panoramic depth features $\mathcal{D}_{t}\in\mathbb{R}^{n\times d_{v}}$ and panoramic RGB features $\mathcal{R}_{t}\in\mathbb{R}^{n\times d_{v}}$, where $d_{v}$ represents the dimension of each view. Moreover, we employ a two-layer transformer-based panorama encoding module to perform self-attention and sequentially encode the depth and RGB panoramic images. The panorama encoding modules for both modalities share weights, since each view in the depth and RGB panoramas is accordant. Subsequently, we aggregate each visited node and its adjacent views by average pooling to construct depth and RGB exploration maps in parallel. The explored depth and RGB trajectory maps are formalized as $\mathcal{T}_{t}^{\mathcal{D}}\in\mathbb{R}^{k\times d_{n}}$ and $\mathcal{T}^{\mathcal{R}}_{t}\in\mathbb{R}^{k\times d_{n}}_{t}$ . Here, $k$ denotes the number of navigable nodes at step $t$, while $d_{n}$ represents the dimension of each navigable node. These exploration maps memorize the agent navigation trajectory information for global action predictions, which serve to guide the agent traversal across various navigable nodes when considering backtracking.
[…]
Figure 4: The pipeline for fusing hybrid environmental representations in the proposed HAP module in Fig. 2.
3.4 Hierarchical Aggregation and Prediction